500cm cube of a saturated solution of ALCL contains 26.6g dissolved in it at 45°c. What is the solubility in moles dm cube at this temperature


1
Expert's answer
2021-06-21T05:30:43-0400

Volume = 500 cm3 which is equal to (500/1000)dm3 = 0.5 dm3

Moles of AlCl3 = 26.6/133.34 = 0.199 moles = 0.2 moles

Solubility = 0.2/0.5 = 0.4 moles/dm3